During the power stroke of an engine, the work produced is 2520 kj/kg and the heat rejected through the cylinder is 570 kj/kg. If the internal energy increased by 120 kj/kg, determine the amount of heat received in kj/s if the working fluid flow at the rate of 3 kg/s.
1
Expert's answer
2020-09-21T08:33:21-0400

According to the first law of thermodynamics, the heat received comes to the increase in the internal energy and work:


"Q=\Delta U+W,\\
Q=120+2520=2640\text{ kJ/kg}."

Also, 570 kJ/kg rejected, which means that heat received was actually greater:


"Q_r=Q+570=3210\text{ kJ/kg}."

Convert this to kJ/s:


"q_r=Q_r\cdot r=3210\text{ kJ/kg}\cdot3\text{ kg/s}=9630\text{ kJ/s}."